[Hallc_running] [New Logentry] Hall C RC summary: Thursday, September 8, 2022

gaskelld at jlab.org gaskelld at jlab.org
Thu Sep 8 08:10:02 EDT 2022

Logentry Text:
Beam returned from the Maintenance Day last night a little after 5 pm, and we have been taking production data on LD2 for the Q=3.85, W=2.62, low epsilon, SHMS left (10 degrees) pi- setting.  We'll switch to dummy at this setting at about 9 am, and then go back to the SHMS center (8 degrees) setting - taking the required dummy data and then going back to LD2 where we will stay until beam goes off tomorrow at 8 am.

The low current BCM calibration originally planned for yesterday has been postponed and is now planned this afternoon (time to be coordinated with Halls B and D since it is invasive).

We will have an RC meeting this morning at 10 am (one hour earlier than our usual meeting time).  The Zoom link is:



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4039117
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220908/3f3b1ded/attachment.html>

More information about the Hallc_running mailing list